OpenStudy (anonymous):

Anyone kind enough to help me out with one question? Will give medal :) 1. Two boys want to use a seesaw, and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

So you are basically asking to make d2 the subject of the formula....right?

OpenStudy (anonymous):

Right

OpenStudy (gorv):

w1*d1=w2*d2 divide both side by w2

OpenStudy (gorv):

\[\frac{ w1*d1 }{ w2 }=\frac{ w2*d2 }{ w2 }\]

OpenStudy (anonymous):

Im confused on that about what the answer will turn out to be after doing that...

OpenStudy (anonymous):

\[\frac{ (d2\times w2) }{ w2 }=\frac{ (w1\times d1) }{ w2 } \therefore. d2=\frac{w1\times d1 }{ w2 }\]

OpenStudy (anonymous):

Okay now so you basically just divide them and get d2=w1xd1 over w2 or is there more steps to this problem? Im kinda getting it and kinda not :/

OpenStudy (anonymous):

Exactly. there aren't a lot of steps involved. Just the dividing

OpenStudy (anonymous):

Oh okay thanks.. you mind helping with another question?

OpenStudy (anonymous):

@Renato19

OpenStudy (anonymous):

2. The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

So in order to make g the subject we have to divide both sides by 0.5t2. Notice that the g is not included because we want it to be alone on the left hand side.  So \[g=\frac{d }{ 0.5t2 }\] Therefore if the basketball takes 10s to hit the ground after being dropped from a height of 490 feet \[g=\frac{ 490 }{ 0.5(10)2 }= 49\]

OpenStudy (anonymous):

Alright thanks :) I really appreciate your help

OpenStudy (anonymous):

Im sorry I have one last question that I have to do and youre really smart and been helping me so can you do one more? its alright if you dont wanna.

OpenStudy (anonymous):

no go ahead

OpenStudy (anonymous):

Okay :) 4. Marcie wants to enclose her yard with a fence. Her yard is in the shape of a triangle attached to a rectangle. See the figure below. The area of this figure can be found by the formula A = (wh) + 0.5(bh). If Marcie wants the total area to be larger than a specified value, she can use the formula A > (wh)+ 0.5(bh). Rewrite this formula to solve for b. Show all steps in your work.

OpenStudy (anonymous):

@Renato19

OpenStudy (anonymous):

OK step 1: take (wh) to the right hand side and change the sign \[A -(wh) > 0.5bh\] Step 2: divide both sides by 0.5h \[\frac{ A-(wh) }{ 0.5h }>\frac{ 0.5bh }{ 0.5h }\] Step 3: The '0.5' and 'h' on the left hand side cancel out leaving only 'b' on the left hand side \[\frac{ A-(wh) }{ 0.5h }>b\]

OpenStudy (anonymous):

Thanks youre the best

OpenStudy (anonymous):

pleasure